# Planning in a recurrent neural network that plays Sokoban

### v2

@@ Appendix C Case Studies
Figure 10: Case studies of three medium-validation levels demonstrating different behaviors after 6 thinking steps. Colors are as in Figure 12. Boxes and targets are paired in upper- and lower-case letters respectively, such that the DRC’s best solution places boxes on targets in alphabetical order. Videos available at this https URL. Levels solved faster obtain higher return as the agent incurs the per-step penalty fewer times. Note that the letters used are for reference and not a property of Sokoban.
@@
For example, thinking lets the DRC solve Figure 10(a). In the no-thinking condition, the DRC first pushes box $C$ one square to the right. It then goes back to push $A$ to $a$, but it has now become impossible to push box $B$ onto $b$. In contrast, after thinking, the DRC pushes $A$ to $a$ first, which lets it solve the level.
@@
Figure 10(b) illustrates a scenario where the agent happens. In the no-thinking condition, the DRC takes many steps back and forth before pushing any boxes. First it goes down to $y$, up to $c$, then down onto $z$, back up to $y$ and to $z$ again. It then proceeds to solve the rest of the puzzle: push box $A$ onto $a$, prepare box $B$ on $x$ and box $C$ where $B$ originally was, push in boxes $B,C$ and finally $D$. In the thinking condition, the DRC makes a beeline for $A$ and then plays the same solution.
@@
Figure 10(c) illustrates a scenario where thinking time results in a slower solution. In the no thinking condition, the DRC starts by pushing box C into position y, then pushes boxes A, then B into place. On the way back down, the DRC pushes C onto c and finally D onto d. In contrast, in the thinking condition, the DRC goes the other way and starts by pushing B onto b. The subsequent solution (A, C, then D) is the same, but the DRC has wasted time trekking back from B to A.
Appendix D Probe Training
@@
We train multiple different linear probes on the hidden states $h$ and cell states $c$ activations of the network. The probes are trained either on the states of the individual layers or by concatenating the states of all the three layers. We collect the activations of the model by letting it play through the hard levels. The training set for the probes are constructed by taking a random sample of cached levels and including all the timesteps in the level excluding the first five timesteps. We exclude these intial steps as we believe they can be noisy as the network comes up with the stategy to solve the levels.
The probes are trained with logistic regression with L1 decay using the Scikit-Learn library. We do a grid-search on the learning rate and the L1 weight decay to select the probe that has the highest F1-score on a validation set.
For multi-class probe targets, each potential output class $l$ is treated like a different data point. That is: for each label $l$, we consider a prediction positive if the highest probe logit is for label $l$, and negative otherwise. We consider a data point positive iff the true label is label $l$. Then, we compute the confusion matrix and the F1 score from the $n\cdot l$ data points.
Appendix E Details about Sparse Autoencoders (SAEs)
@@
In order to search for more monosemantic and interpretable features in the network, we train sparse-autoencoders (SAEs) [Huben et al., 2023, Bricken et al., 2023] on the individual squares in the $h$ hidden state of the network consisting of 32 neurons. Thus, we get a $10\times 10$ visualization for each SAE feature as shown in Figure 11. We use the top-k activation function [Gao et al., 2024, Tamkin et al., 2024] for the SAE as it is currently the state-of-the-art method to train SAEs and is easier to work with as it directly sets the $L_{0}$ norm of the SAE activations. The hyperparameter search space for training the SAE is provided in Table 7. We train separate SAEs for each layer with the specified hyperparameters and pick the one that achieves greater than $90\%$ explained variance while having interpretable features assessed through manual visual inspection. We release these trained SAEs and probes in the same huggingface repo as our trained DRC networks.  [^8] https://huggingface.co/AlignmentResearch/learned-planner
We provide examples of some of the interpretable features from an SAE trained on the last layer with $k=8$ in Table 8 with corresponding visualization in Figure 11. For the “Target”, “Unsolved”, and “Solved” concepts (Table 10), we observed cases where the SAE feature is offset from the ground truth by 1 square in either horizontal or vertical directions. The level setup includes a permanent one-square outer-edge wall, so this offset never results in an out-of-bounds issue. We evaluated these potential “Offset” variants for the Target, Unsolved, and Solved concepts.
All of the features in the SAE that we find to be interpretable are already embedded in individual channels. These channels are either as monosemantic as the SAE features, or in some cases more monosemantic as measured by the F1 score. Table 3 reports the precision, recall, and F1 scores for the action features as evaluated through channels, SAE neurons, and linear probe trained against the ground truth action predictions. Table 10 reports the same scores for additional interpretable features. The SAE action features have lower F1 scores than channels by a margin of $5.9\%$ on average. On the other hand, the linear probes trained across all channels of the hidden state have similar F1 scores as the specified channels indicating that the channels are monosemantic and cannot be improved upon by combining various channels using a linear probe.
@@
Appendix F Additional quantitative behavior figures and tables
Figure 12: Left: A Sokoban level from the hard set. Right: the same level as the NN sees it, one pixel per tile. Walls are black, boxes are brown, targets are pink and the robot is green.
Figure 13: Number of thinking steps required to solve the level vs. number of nodes A* needs to expand to solve it. The trend is increasing at the end but very unclear, indicating different heuristics used by the DRC and A*.
Figure 14: Success rate on datasets of various difficulty, for various checkpoints of each architecture. This deviates very little from a curve, which shows that ResNets and DRCs which are equally good at the easier sets are also equally good at the harder sets. Perhaps DRC(1,1) is a slight exception, but it also performs much worse than the others overall (see Figure 7).
@@ Appendix F Additional quantitative behavior figures and tables
(c) with cycles from those levels that were solved by the agent
Figure 15: We replace $N$-length cycles with $N$ thinking steps and check for the same state after some timesteps. (a) A histogram of cycle lengths in the medium-validation set. (b, c) After replacing a cycle with the same length in thinking steps, are all the states the same for the next $x$ steps?

### v3

@@ Appendix E Case Studies
Figure 18: Case studies of three medium-validation levels demonstrating different behaviors after 6 thinking steps. Colors are as in Figure 1 (right). Boxes and targets are paired in upper- and lower-case letters respectively, and the optimal solution places boxes alphabetically. Videos available at this https URL. Levels solved faster incur fewer per-step penalties, yielding higher returns. Letters are for reference only and not intrinsic to Sokoban.
@@
In this scenario, the DRC fails to solve the level in the no-thinking condition. It initially pushes box $C$ one square to the right. While attempting to push $A$ to $a$, it blocks $B$ from reaching $b$, making the level unsolvable. In the thinking condition, the DRC correctly pushes $A$ to $a$ first, leaving room to push $B$ to $b$ and $C$ to ts target, solving the level.
@@
Without thinking, the DRC takes inefficient route, moving back and forth multiple times before starting to push boxes. For example, it moves to $y$, up to $c$, then down onto $z$, back up to $y$ and to $z$ before solving the puzzle. With thinking, the DRC immediately moves to box $A$, completing the solution with fewer extraneous steps. The resulting solution is faster, reducing the total step penalty.
@@
In this level, thinking causes the DRC to take a less efficient path. Without thinking, the DRC pushes box $C$ to $y$, followed by boxes $A$, then $B$ into place. On the way back down, the DRC pushes $C$ onto $c$ and finally $D$ onto $d$. In the thinking condition, the DRC starts by pushing B onto b then backtracks to push A, C, then D. This unnecessary backtracking leads to a slower overall solution.
Appendix F Probe Training
@@
We train multiple linear probes on the hidden states $h$ and cell states $c$ activations of the network. Probes are trained either on activations from individual layers or by concatenating activations across all three layers. Activations were collected as the model played through hard levels. The training set are constructed by randomly sampling cached levels and including all timesteps except the first five. These initial steps excluded due to potential noise while the network formulated its strategy for solving the levels.
Probes are trained with logistic regression with L1 decay using the Scikit-Learn library. A grid search over learning rates and L1 weight decay was conducted to identify the probe with the highest F1 score on the validation set.
For multi-class probe targets, each potential output class $l$ is treated as a separate data point. Specificially: A prediction was considered positive if the highest probe logit corresponded to class $l$, and negative otherwise. A data point was positive if its true label matched $l$. The confusion matrix and F1 score were then computed from the $n\cdot l$ data points.
In addition to Agent-Directions and Box-Directions probes, we trained the following:
  • Next-Action probe: Grid-wise binary probe for each of the four move action that predict $1$ for all squares in the grid if the next action matches the probe’s action. We train grid-wise probe instead of a global probe so that the probe can be applied to arbitrary-sized inputs. The results for the next-action probes are available in Table 10.
Probe results.
The pacing probe gets $F_{1}=31.0\%$, which is not much better than the constant 1 probe, which has $F_{1}=12.8\%$. This suggests that the $\text{DRC}(3,3)$ does not represent whether it is in a cycle or not.
The possible box directions probe got approximately $\approx 30\%$ accuracy, and it was
Table 5: Weight and bias for transforming action probes to predictions
Mean Max Positive proportion
Weight 1.2086 -0.0582 0.2070
Up Down Left Right
Bias 0.3337 -0.0921 -0.0632 -0.0539
Appendix G Looking for interpretable features with Sparse Autoencoders (SAEs)
@@
To search for monosemantic and interpretable features in the network, we train sparse-autoencoders (SAEs) [Huben et al., 2023, Bricken et al., 2023] on the individual squares in the $h$ hidden state of the network consisting of 32 neurons. Thus, we get a $10\times 10$ visualization for each SAE feature as shown in Figure 19. We use the top-k activation function [Gao et al., 2024, Tamkin et al., 2024], which is state-of-the-art for training SAEs and directly enforces an $L_{0}$ sparcity constraint on activations. The SAE hyperparameter search space is detailed in Table 6. We train separate SAEs for each layer with the specified hyperparameters, selecting those achieving greater than $90\%$ explained variance while retaining interpretable features based on manual visual inspection. We release these trained SAEs and probes in the huggingface repository with our trained DRC networks.  [^8] https://huggingface.co/AlignmentResearch/learned-planner
Examples of interpretable features from an SAE trained on the last layer with $k=8$ are provided in Table 7, with visualizations in Figure 19. For “Target”, “Unsolved”, and “Solved” concepts (Table 9), we occasionally observed an offset of 1 square (horizontal or vertical) from the ground truth. Due to a permanent one-square outer-edge wall, this offset never results in out-of-bounds errors. We evaluated these potential “Offset” variants for the Target, Unsolved, and Solved concepts.
All interpretable SAE features identified are already embedded in individual channels, with comparable or higher monosemanticity then the SAE features based on F1 scores. Table 10 compares precision, recall, and F1 scores for action features across channels, SAE neurons, and linear probes trained against the ground truth predictions. Table 9 reports scores for other interpretable features. SAE action features consistently underperform channels by an average F1 margin of $5.9\%$. Linear probes trained across all hidden-state channels achieve similar F1 scores to individual channels indicating the inherent monosemanticity of channels without further improvement through linear probe.
@@
Feature Channel SAE Feature Probe
Number Prec Rec F1 Number Prec Rec F1 Prec Rec F1
Up L3C29 95.7 88.1 91.7 L3F270 93.9 76.2 84.1 97.5 86.5 91.7
Down L3C8 98.4 80.8 88.8 L3F187 98.0 79.1 87.6 97.6 86.9 91.9
Left L3C27 85.5 84.6 85.1 L3F244 96.1 63.2 76.2 83.5 86.6 85.0
Right L3C3 97.0 86.9 91.7 L3F385 94.6 78.5 85.8 97.6 87.4 92.2
Appendix H Additional quantitative behavior figures and tables
Figure 21 through Figure 23 present quantitative behavior, analyzing various aspects of the network’s planning behavior and evaluation metrics.
Figure 20: Number of thinking steps required to solve the level vs. number of nodes A* needs to expand to solve it. The weak correlation toward the end indicates different that the DRC and A* rely on different heuristics.
Figure 21: Success rate on datasets of varying difficulty for different architecture checkpoints. Performance trends suggest ResNets and DRCs with comparable results on easier sets also perform similarly on harder sets. DRC(1,1) is a slight exception, performing consistently worse overall (see Figure 8).
@@ Appendix H Additional quantitative behavior figures and tables
(c) with cycles from levels solved by agent
Figure 22: We replace $N$-length cycles with $N$ thinking steps to examine state consistncy across subsequent timesteps. (a) Histogram of cycle lengths in the medium-validation set. (b, c) After replacing a cycle with the same length in thinking steps, are all the states the same for the next $x$ steps?
(a) Pacing behavior on file 0, level 53. On the given starting observation, the agent paces around 4 spaces in the first 9 steps and then goes on to solve the level. Video for the level is available at this https url .
(b) Change in per-step F1 score of Box-Directions probe for moves in cycles and outside cycles on medium-difficulty validation levels. The non-cycle moves were recorded from the same distribution of timesteps where cycles occur but from levels without a cycle at those steps. Mean per-step change in F1 for cycle and non-cycle steps are $1.40\%\pm 0.06\%$ and $0.84\%\pm 0.04\%$ respectively.
Figure 23: Illustration of cycles and F1 scores
Appendix I Additional related work
Ethical treatment of AIs.
The question whether AIs deserve moral consideration has been widely debated. Schwitzgebel and Garza [2015] argue that highly human-like AIs merit rights, while Tomasik [2015] suggest that most AIs deserve some degree of consideration, akin to biological organisms [Singer, 2004]. The concept of ethical treatement for reinforcement learners has been explored by Daswani and Leike [2015], who propose that pleasure and pain in these systems may correspond to temporal difference (TD) error rather than absolute returns. If neural networks have internal objectives distinct from their critic head [Hubinger et al., 2019, Di Langosco et al., 2022], identifying these objectives could provide higher-assurance methods for assessing AI goals compared to direct querying [Perez and Long, 2023].
Chain-of-thought faithfulness.
The faithfulness of chain of thought reasoning in large language models (LLMs). Studies such as [Lanham et al., 2023, Pfau et al., 2024] investigate whether LLMs rely on plain English for long-term reasoning, which could allow unintended consequences to be easily identified and mitigated Scheurer et al. [2023].
Fully reverse engineering small networks.
Recent efforts have successfully reverse-engineered small neural networks performing algorithmic tasks [Nanda et al., 2023a, Chughtai et al., 2023, Zhong et al., 2023, Quirke and Barez, 2023].
Systematic Generalization.
Previous work identified conditions under which neural networks generalize, such as diverse datapoints and egocentric environments [Lake and Baroni, 2023, Hill et al., 2020, Mutti et al., 2022]. Similar interpretability can be extended across these neural networks to uncover shared planning mechanisms and the conditions in which they emerge.
Appendix J Sussman’s anomaly
Although the $\text{DRC}(3,3)$ shows exceptional long-term planning capabilities as demonstrated in the paper, it can have some trivial failure modes. Figure 24 shows a level in which the agent indefinitely tries to put the two boxes on the right onto the same target that is on the right side of the level. This failure mode is similar to the Sussman’s anomaly demonstrated by Sussman [1973] illustrating the weakness of non-interleaved planning algorithms. We observe that such cases happen rarely, with the network being able to resolve the interleaved dependencies between boxes in most cases.
Figure 24: Level 466 from file 233 of train-medium in which the $\text{DRC}(3,3)$ network tries to put both the boxes on the right to the target on the right by repeatedly putting one box on target, only to remove it and put the other one on the same target.
Appendix K Planning vs. predicting box-directions
Most of our analysis and experimentation in the paper relies on the box-directions probe, which predicts the sequence of all future box movement directions from squares in the grid given network’s activations. While predicting the box-directions is very close to planning, it misses a few details required for a complete plan. For example, for a complete plan of actions, the agent also has to also know the order in which it wants to move the boxes in and the plan a path from the agent’s current position to the next box to push. However, we do train probes for agent-directions and next-box to move and show they have high predictive accuracy. We focus primarily on boxes-directions because it is the most crucial component of a plan in the game, which is also reflected in the fact that the box-directions probe is the most causal probe.